OpenStudy (anonymous):

49.6=2(9.7)+2(s) solve for s

jimthompson5910 (jim_thompson5910):

49.6=2(9.7)+2s 49.6=19.4+2s 49.6-19.4 = 2s 30.2 = 2s 30.2/2 = s 15.1 = s s = 15.1